Use the compound interest formula to determine the interest earned in the given period.
p=$4000 at 6.5% compounded annually for 8 years.
The previous answer I was given was wrong.
The correct answer is $2619.98 but how do I get that?
Thanks!

Answer by Theo(13342)

4000 * (1.065)^8 = 6619.982685
6619.982685 - 4000 = 2619.982685 which rounds to 2619.98.

the formula required is f = p * (1+r)^n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

6.5% annual percentage rate compounded annually is equal to .065.
r = .065

n is the number of time periods.
since the time period is in years, then n = 8 years.

the interest earned is equal to the future value minus the present value.
if i is the the interest, then i = f - p.